# Extreme Compression of Large Language Models via Additive Quantization

### v2

@@ Abstract
The emergence of accurate open large language models (LLMs) has led to a race towards performant quantization techniques which can enable their execution on end-user devices. In this paper, we revisit the problem of “extreme” LLM compression—defined as targeting extremely low bit counts, such as 2 to 3 bits per parameter—from the point of view of classic methods in Multi-Codebook Quantization (MCQ). Our algorithm, called AQLM, generalizes the classic Additive Quantization (AQ) approach for information retrieval to advance the state-of-the-art in LLM compression, via two innovations: 1) learned additive quantization of weight matrices in input-adaptive fashion, and 2) joint optimization of codebook parameters across entire layer blocks. Broadly, AQLM is the first scheme that is Pareto optimal in terms of accuracy-vs-model-size when compressing to less than 3 bits per parameter, and significantly improves upon all known schemes in the extreme compression (2bit) regime. In addition, AQLM is practical: we provide fast GPU and CPU implementations of AQLM for token generation, which enable us to match or outperform optimized FP16 implementations for speed, while executing in a much smaller memory footprint.
@@ 1 Introduction
    We propose the AQLM algorithm, which extends AQ to post-training compression of LLM weights, via two innovations: (1) adapting the MAP-MRF optimization problem behind AQ to be instance-aware, taking layer calibration input & output activations into account; (2) complementing the layer-wise optimization with an efficient intra-layer tuning technique, which optimizes quantization parameters jointly over several layers, using only the calibration data.
@@
    We evaluate the effectiveness of this algorithm on the task of compressing accurate open LLMs from the Llama 2 (Touvron et al., 2023) family with compression rates of 2-4 bits per parameter. We find that AQLM outperforms the previous state-of-the-art across the standard 2-4 bit compression range, with the most significant improvements for extreme 2-bit quantization (see Figure 1). We provide detailed ablations for the impact of various algorithm parameters, such as code width and number of codebooks, and extend our analysis to the recent Mixtral model (Jiang et al., 2024).
@@
    We show that AQLM is practical, by providing efficient GPU and CPU kernels implementations for specific encodings, as well as end-to-end generation [^1] https://github.com/vahe1994/AQLM. Results show show that our approach can match or even outperform the floating point baseline in terms of speed, while reducing the memory footprint by up to 8x. Specifically, AQLM can be executed with layer-wise speedups of $\sim 30\%$ for GPUs, and of up to 4x for CPU inference.
@@ 2.1 LLM Quantization
The state-of-the-art method in terms of accuracy-to-size trade-off is QuIP (Chee et al., 2023). Concurrent to our work, an improved variant called QuIP# (Tseng et al., 2023) was introduced. Roughly, these methods work by first “smoothening” weights by multiplying with a rotation matrix, and then mapping them onto a lattice. QuIP was the first method to obtain stable results (i.e., single-digit PPL increases) in the 2-bit per parameter compression range. At a high level, QuIP and QuIP# aim to minimize the “worst-case” error for each layer, given initial weights and calibration data. For instance, in QuIP#, the distribution of the rotated weights approximates a Gaussian, while the encoding lattice (E8P) is chosen to minimize “rounding” error. By contrast, our approach uses a different weight encoding (codebooks are additive), and learned codebooks instead of a fixed codebook. Thus, our insight is that we should be able to obtain higher accuracy by direct optimization of the codebooks over the calibration set, removing the rotation. Further, we show that codebooks for different layers can co-train via joint fine-tuning over the calibration data.
@@ 2.2 Quantization for Nearest Neighbor Search
Finally, several recent works (Martinez et al., 2016, 2018; Zhang et al., 2014) elaborate the idea of Additive Quantization, proposing the more effective procedure for codebooks learning. Composite Quantization (CQ) (Zhang et al., 2014) learns codebooks with a fixed value of scalar product between the codewords from different codebooks. Currently, the state-of-the-art compression accuracy is achieved by the LSQ method (Martinez et al., 2018).
@@ 3.1 Overview
In the following, we will assume that $\widehat{\mathbf{W}}$ is quantized using AQ, and adopt standard notation (Martinez et al., 2016). AQ splits weight rows into groups of $g$ consecutive elements, and represents each group of weights as a sum of $M$ vectors chosen from multiple learned codebooks $C_{1},...C_{M}$, each containing $2^{B}$ vectors (for B-bit codes). A weight is encoded by choosing a single code from each codebook and summing them up. We denote this choice as a one-hot vector $b_{m}$, which results in the following representation for a group: $\sum_{m=1}^{M}C_{m}b_{ijm}$. This is similar to PTQ algorithms (Frantar et al., 2022a), except for using much more complex coding per group. To represent the full weights, we simply concatenate:
@@
To learn this weight representation, we initialize codebooks $C$ and codes $b$ by running residual K-means as in Chen et al. (2010). Then, we alternate between updating codes $b_{i,j,m}$ and codebooks $C_{m}$ until the loss function (3) stops improving up to the specified tolerance. Since codes are discrete and codebooks are continuous, and we are optimizing over multiple interacting layers, our approach has three phases, described in Algorithm 1 and detailed below.
@@ 3.2 Phase 1: Beam search for codes
Above, $\langle\cdot,\cdot\rangle_{F}$ denotes a Frobenius inner product of two matrices. Next, let us consider the three components of Eqn. (4) in isolation. First, note that $||\mathbf{W}\mathbf{X}||_{2}^{2}$ is constant in $b$ and can be ignored. The second component can be expanded further into pairwise dot products:
@@ 3.3 Phase 2: Codebook update
For each update phase, our implementation runs 100 Adam steps with learning rate 1e-4. However, we found that the final result is not sensitive to either of these parameters: training with smaller number of steps or learning rate achieves the same loss, but takes longer to converge. In future work, these hyperparameters could be eliminated by switching to dedicated least squares solver for codebooks. Similarly to other algorithms, we also learn per-unit scales $s\in\mathbb{R}^{h}$ that are initialized as $s_{i}:=||\mathbf{W}_{i}||_{2}$ and updated alongside codebooks via the same optimizer (line 19 in Algorithm 1).
@@ 3.4 Phase 3: Fine-tuning for intra-layer cohesion
Here, we opt for a middle ground by performing optimization at the level of individual transformer blocks, i.e. groups of 4-8 linear layers [^2] This number depends on factors including the use of gated GLU activations, group query attention and QKA weight merging. that constitute a single multi-head self-attention, followed by a single and MLP layer. Having quantized all linear layers within a single transformer block, we fine-tune its remaining parameters to better approximate the original outputs of that transformer block by backpropagating through the weight representation (2).
@@
Table 1: Evaluation of quantized Llama 2 models for 2-2.8 bits per parameter, with an extra section for higher bitwidth. We report perplexity on WikiText2 (Merity et al., 2016) & C4 (Raffel et al., 2020) and accuracy for zero-shot tasks. The Average accuracy is the mean of 5 zero-shot tasks. Primary metrics are Wiki2 (PPL), C4 (PPL) and Average accuracy.
Size Method Avg bits Wiki2normal-$\downarrow$ C4normal-$\downarrow$ WinoGrandenormal-$\uparrow$ PiQAnormal-$\uparrow$ HellaSwagnormal-$\uparrow$ ArcEnormal-$\uparrow$ ArcCnormal-$\uparrow$ Average accuracynormal-$\uparrow$
@@
Size Method Avg bits Wiki2normal-$\downarrow$ C4normal-$\downarrow$ WinoGrandenormal-$\uparrow$ PiQAnormal-$\uparrow$ HellaSwagnormal-$\uparrow$ ArcEnormal-$\uparrow$ ArcCnormal-$\uparrow$ Average accuracynormal-$\uparrow$
@@ 4.1 Compression quality for modern LLMs
We report perplexity on WikiText2 (Merity et al., 2016) and C4 (Raffel et al., 2020) validation sets. We also measure zero-shot accuracy on WinoGrande (Sakaguchi et al., 2021), PiQA (Tata & Patel, 2003), HellaSwag (Zellers et al., 2019), ARC-easy and ARC-challenge (Clark et al., 2018) via the LM Eval Harness (Gao et al., 2021). We broadly follow the evaluation setup of GPTQ (Frantar et al., 2022a).
We consider three main targets in terms of compression ranges: 2-2.8 bits, 3-3.1 bits, and 4-4.1 bits per model parameter. In the results below average bits per parameter takes into account only quantized weights, we do not include parameters kept in floating precision similarly to the related work. The details on the model size estimate are provided in Appendix G. We compare AQ against GPTQ for 3&4 bits (Frantar et al., 2022a), SpQR for 3&4 bits (Dettmers et al., 2023b), QuIP in 2,3 & 4 bits (Chee et al., 2023) and QuIP# for 2&4 bits (Tseng et al., 2023). While GPTQ and SpQR technically support 2-bit quantization, they perform poorly in the 2-3 bit range. For QuIP, we omit results for the 7B model, as we could not achieve competitive performance in this one scenario using the available implementations. (Currently, there is no official implementation of the original QuIP (non-#) for the Llama 2 model.) For QuIP#, we focus on 2 and 4 bit because the available implementation does not yet support 3-bit compression. We calibrate each algorithm using the subset of RedPajama dataset (Computer, 2023), with a sequence length of 4096.
The exact bit-widths for each method are dictated by parameters such as the number of codebooks and code width. We report results for the $2{-}2.8$ and $3{-}3.1$ bitwidth ranges in Tables 1 and 2, respectively. Additional results for $4-4.1$ bits are deferred to Appendix E.2.
Table 3: Evaluation of quantized Mixtral (Jiang et al., 2024) models for 2 bits. The table reports perplexity on WikiText2 (Merity et al., 2016) and C4 (Raffel et al., 2020), as well as accuracy for zero-shot tasks. The Average accuracy column is the mean of 5 zero-shot task accuracies. Primary metrics are Wiki2 (PPL), C4 (PPL) and Average accuracy.
Size Method Avg bits Wiki2normal-$\downarrow$ C4normal-$\downarrow$ WinoGrandenormal-$\uparrow$ PiQAnormal-$\uparrow$ HellaSwagnormal-$\uparrow$ ArcEnormal-$\uparrow$ ArcCnormal-$\uparrow$ Average accuracynormal-$\uparrow$
@@
Table 3 presents results on the Mixtral MoE-type model, comparing against QuIP# at 2-bits. (See Appendix E.1 for full results.) AQLM outperforms QuIP# in this case as well. Although the margins are lower compared to Llama 2 models, they are still significant for “harder” tasks, such as Arc Challenge (+3 points).
@@
AQLM compression to strictly 2 bits for the same model is also below Pareto-optimality, as it is outperformed by 4-bit AQLM compression for Llama 2 7B (5.21 vs 5.65). To find the Pareto-optimal quantization bitwidth, we run experiments between 2-3 bits per parameter and report them in Table 1, below horizontal bars. Thus, the Pareto-optimal bitwidth for AQLM appears to be around 2.5 bits per parameter (Table 1), at which point we are comparable to 5-bit AQLM for Llama 2 7B (Appendix Table 10). In turn, the 2.76-bit AQLM on 13B outperforms the uncompressed 7B model. As such, AQLM is the first algorithm to achieve Pareto-optimality at less than 3 bits per parameter.
@@ 4.2 Ablation analysis
In Appendix D, we examine key design choices regarding initialization, alternating optimization, the impact of the fine-tuning protocol, distribution of codebooks vs groups, as well as other hyper-parameters. In brief, we first find that the residual K-means initialization is critical for fast algorithm convergence: when compared with random initialization, it needs significantly fewer training iterations. Second, to validate our calibration fine-tuning procedure, we compare it against 1) no fine-tuning, 2) fine-tuning only of non-linear layers (e.g. RMSNorm) but not of codebook parameters, and 3) fine-tuning only the codebooks (but not other layers). The results, presented in full in Appendix D, show that fine-tuning the codebook parameters has the highest impact on accuracy, by far, while fine-tuning the RMSNorm only has minor impact. This validates our choice of leveraging the calibration set for learned codebooks.
Further, we observe that increasing the number of sample sequences in the range 128 to 4096 leads to a gradual PPL improvement, but with diminishing returns. In this respect, AQLM benefits more from larger calibration sets (similarly to QuIP#), as opposed to direct methods like GPTQ which saturate accuracy at around 256 input sequences. Finally, we investigate various options for investing a given bit budget, comparing e.g. longer codes (e.g. 1x15) vs multiple codebooks with shorter codes (e.g. 2x8).

### v3

@@ Abstract
The emergence of accurate open large language models (LLMs) has led to a race towards performant quantization techniques which can enable their execution on end-user devices. In this paper, we revisit the problem of “extreme” LLM compression—defined as targeting extremely low bit counts, such as 2 to 3 bits per parameter—from the point of view of classic methods in Multi-Codebook Quantization (MCQ). Our algorithm, called AQLM, generalizes the classic Additive Quantization (AQ) approach for information retrieval to advance the state-of-the-art in LLM compression, via two innovations: 1) learned additive quantization of weight matrices in input-adaptive fashion, and 2) joint optimization of codebook parameters across each transformer blocks. Broadly, AQLM is the first scheme that is Pareto optimal in terms of accuracy-vs-model-size when compressing to less than 3 bits per parameter, and significantly improves upon all known schemes in the extreme compression (2bit) regime. In addition, AQLM is practical: we provide fast GPU and CPU implementations of AQLM for token generation, which enable us to match or outperform optimized FP16 implementations for speed, while executing in a much smaller memory footprint.
@@ 1 Introduction
    We propose the AQLM algorithm, which extends AQ to post-training compression of LLM weights, via two innovations: (1) adapting the MAP-MRF [^1] Maximum a Posteriori inference in Markov Random Fields optimization problem behind AQ to be instance-aware, taking layer calibration input & output activations into account; (2) complementing the layer-wise optimization with an efficient intra-block tuning technique, which optimizes quantization parameters jointly over several layers, using only the calibration data.
@@
    We evaluate the effectiveness of this algorithm on the task of compressing accurate open LLMs from the Llama 2 (Touvron et al., 2023) family with compression rates of 2-4 bits per parameter. We find that AQLM outperforms the previous state-of-the-art across the standard 2-4 bit compression range, with the most significant improvements for extreme 2-bit quantization (see Figure 1). We provide detailed ablations for the impact of various algorithm parameters, such as code width and number of codebooks, and extend our analysis to the recent Mixtral model (Jiang et al., 2024). We also evaluate AQLM with improved fine-tuning algorithms from subsequent works, which leads to further increase in accuracy for 2- and 3-bit models.
@@
    We show that AQLM is practical, by providing efficient GPU and CPU kernels implementations for specific encodings, as well as end-to-end generation [^2] https://github.com/Vahe1994/AQLM. Results show that our approach can match or even outperform the floating point baseline in terms of speed, while reducing the memory footprint by up to 8x. Specifically, AQLM can be executed with layer-wise speedups of $\sim 30\%$ for GPUs, and of up to 4x for CPU inference.
@@ 2.1 LLM Quantization
The published state-of-the-art method is QuIP (Chee et al., 2023). Concurrent to our work, an improved variant called QuIP# (Tseng et al., 2024) was introduced. Roughly, they work by first “smoothening” weights by multiplying with a rotation matrix, and then mapping them onto a lattice. At a high level, QuIP and QuIP# aim to minimize the “worst-case” error for each layer, given initial weights and calibration data. For instance, in QuIP#, the distribution of the rotated weights approximates a Gaussian, while the encoding lattice (E8P) is chosen to minimize “rounding” error. By contrast, our approach uses a different weight encoding (codebooks are additive), and learned codebooks instead of a fixed codebook. Thus, our insight is that we should be able to obtain higher accuracy by direct optimization of the codebooks over the calibration set, removing the rotation. Further, we show that codebooks for different layers can co-train via joint fine-tuning over the calibration data.
@@ 2.2 Quantization for Nearest Neighbor Search
Finally, several recent works (Martinez et al., 2016, 2018; Zhang et al., 2014) elaborate the idea of Additive Quantization, proposing the more effective procedure for codebooks learning. Composite Quantization (CQ) (Zhang et al., 2014) learns codebooks with a fixed value of inner product between the codewords from different codebooks. Currently, the state-of-the-art compression accuracy is achieved by the LSQ method (Martinez et al., 2018).
@@ 3.1 Overview
In the following, we will assume that $\widehat{\mathbf{W}}$ is quantized using AQ, and adopt standard notation (Martinez et al., 2016). AQ splits weight rows into groups of $g$ consecutive elements, and represents each group of weights as a sum of $M$ vectors chosen from multiple learned codebooks $C_{1},...,C_{M}$, each containing $2^{B}$ vectors (for $B$-bit codes). A weight is encoded by choosing a single code from each codebook and summing them up. We denote this choice as a one-hot vector $b_{m}$, which results in the following representation for a group: $\sum_{m=1}^{M}C_{m}b_{ijm}$. This is similar to PTQ algorithms (Frantar et al., 2022a), except for using much more complex coding per group. To represent the full weights, we simply concatenate:
@@
To learn this weight representation, we initialize codebooks $C$ and codes $b$ by running residual K-means as in Chen et al. (2010). Specifically, the initialization algorithm proceeds as follows: first, it runs K-means clustering of weight groups and saves the resulting cluster indices. Next, it computes the quantization errors by subtracting the nearest cluster from every weight. Finally, the algorithm runs another round of K-means clustering, but this time on quantization errors instead of weights. Thus, each subsequent codebook is initialized to compensate the quantization error from previous codebooks. After initialization, we alterQuantizingnate between updating codes $b_{i,j,m}$ and codebooks $C_{m}$ until the loss function (3) stops improving up to the specified tolerance. Since codes are discrete and codebooks are continuous, and we are optimizing over multiple interacting layers, our approach has three phases, described in Algorithm 1 and detailed below.
@@ 3.2 Phase 1: Beam search for codes
Above, $\langle\cdot,\cdot\rangle_{F}$ denotes a Frobenius inner product of two matrices. Next, let us consider the three components of Eqn. (4) in isolation. First, note that $||\mathbf{W}\mathbf{X}||_{2}^{2}$ is constant in $b$ and can be ignored. The third component can be expanded further into pairwise dot products:
@@ 3.3 Phase 2: Codebook update
For each update phase, our implementation runs 100 Adam steps with learning rate $10^{-4}$. However, we found that the final result is not sensitive to either of these parameters: training with smaller number of steps or learning rate achieves the same loss, but takes longer to converge. In future work, these hyperparameters could be eliminated by switching to dedicated least squares solver for codebooks. Similarly to other algorithms, we also learn per-unit scales $s\in\mathbb{R}^{d_{out}}$ that are initialized as $s_{i}:=||\mathbf{W}_{i}||_{2}$ and updated alongside codebooks via the same optimizer (line 10 in Algorithm 1).
@@ 3.4 Phase 3: Fine-tuning for intra-layer cohesion
Here, we opt for a middle ground by performing optimization at the level of individual transformer blocks, i.e. groups of 4-8 linear layers [^3] This number depends on factors including the use of gated GLU activations, group query attention and QKV weight merging. that constitute a single multi-head self-attention, followed by a single MLP layer. Having quantized all linear layers within a single transformer block, we fine-tune its remaining parameters to better approximate the original outputs of that transformer block by backpropagating through the weight representation (2).
@@
Table 1: Evaluation of quantized Llama 2 models for 2-2.8 bits per parameter, with an extra section for higher bitwidth. We report perplexity on WikiText-2 (Merity et al., 2016) & C4 (Raffel et al., 2020) and accuracy for zero-shot tasks. The Average accuracy is the mean of 5 zero-shot tasks. Primary metrics are Wiki2 (PPL), C4 (PPL) and Average accuracy.
[…]
@@
[…]
@@ 4.1 Compression quality for modern LLMs
We report perplexity on WikiText-2 (Merity et al., 2016) and C4 (Raffel et al., 2020) validation sets. We also measure zero-shot accuracy on WinoGrande (Sakaguchi et al., 2021), PiQA (Tata & Patel, 2003), HellaSwag (Zellers et al., 2019), ARC-easy and ARC-challenge (Clark et al., 2018) via the LM Eval Harness (Gao et al., 2021). We follow the evaluation setup of GPTQ (Frantar et al., 2022a) and provide configurations for AQLM and baselines in Appendix C.
We consider three main targets in terms of compression ranges: 2-2.8 bits, 3-3.1 bits, and 4-4.1 bits per model parameter. In the results below average bits per parameter takes into account only quantized weights, we do not include parameters kept in floating precision similarly to the related work. The details on the model size estimate are provided in Appendix H. We compare AQ against GPTQ for 3&4 bits (Frantar et al., 2022a), SpQR for 3&4 bits (Dettmers et al., 2023b), QuIP in 2,3 & 4 bits (Chee et al., 2023) and QuIP# for 2&4 bits (Tseng et al., 2024). While GPTQ and SpQR technically support 2-bit quantization, they perform poorly in the 2-3 bit range. For QuIP, our adapted [^4] The official QuIP (non-#) code does not support Llama 2. implementation shows acceptable performance for Llama 2 13B & 70B but performs poorly for the 7B model. We calibrate each algorithm using the subset of RedPajama dataset (Computer, 2023), with a sequence length of 4096.
Table 3: Evaluation of quantized Mixtral (Jiang et al., 2024) models for 2 bits. The table reports perplexity on WikiText-2 (Merity et al., 2016) and C4 (Raffel et al., 2020), as well as accuracy for zero-shot tasks. The Average accuracy column is the mean of 5 zero-shot task accuracies. Primary metrics are Wiki2 (PPL), C4 (PPL) and Average accuracy.
[…]
@@
The exact bit-widths for each method are dictated by parameters such as the number of codebooks and code width. We report results for the $2{-}2.8$ and $3{-}3.1$ bitwidth ranges in Tables 1 and 2, respectively. Additional results for $4-4.1$ bits are deferred to Appendix F.2.
@@
Table 3 presents results on the Mixtral MoE, comparing against QuIP# at 2-bits. (See Appendix F.1 for full results.) AQLM outperforms QuIP# in this case as well. Although the margins are lower compared to Llama 2 models, they are still significant for “harder” tasks, such as Arc Challenge (+3 points).
@@
AQLM compression to strictly 2 bits for the same model is also below Pareto-optimality, as it is outperformed by 4-bit AQLM compression for Llama 2 7B (5.21 vs 5.59). To find the Pareto-optimal quantization bitwidth, we run experiments between 2-3 bits per parameter and report them in Table 1, below horizontal bars. Thus, the Pareto-optimal bitwidth for AQLM appears to be around 2.5 bits per parameter (Table 1), at which point we are comparable to 5-bit AQLM for Llama 2 7B (Appendix Table 10). In turn, the 2.76-bit AQLM on 13B outperforms the uncompressed 7B model. As such, AQLM is the first algorithm to achieve Pareto-optimality at less than 3 bits per parameter.
Table 4: Evaluation of quantized Llama 2 with end-to-end fine-tuning, with the same metrics as in Table 1.
[…]
7B 16 5.12 6.63 67.25 78.45 56.69 69.32 40.02 62.35
AQLM 2.02 6.14 8.09 65.67 76.01 51.83 63.43 34.39 58.27
QuIP# 2.02 6.19 8.16 64.96 75.41 51.91 64.96 35.15 58.48
AQLM 2.29 5.92 7.86 63.77 76.93 52.82 66.16 36.95 59.33
13B 16 4.57 6.05 69.61 78.73 59.72 73.27 45.56 65.38
AQLM 1.97 5.33 7.19 68.67 76.82 56.31 69.99 40.36 62.43
QuIP# 2.01 5.35 7.20 67.64 77.26 56.04 69.02 39.85 61.96
AQLM 2.19 5.22 6.98 68.27 77.53 57.09 69.78 40.70 62.67
70B 16 3.12 4.97 76.95 81.07 63.99 77.74 51.11 70.17
AQLM 2.07 3.83 5.62 74.35 80.90 62.17 74.58 48.98 68.20
QuIP# 2.01 3.91 5.71 74.66 79.54 62.52 77.06 47.61 68.28
4.2 End-to-end fine-tuning experiments
[…]
@@ 4.3 Ablation analysis
In Appendix E, we examine key design choices regarding initialization, alternating optimization, the impact of the fine-tuning, and sensitivity to hyperparameters. In brief, we first find that the residual K-means initialization is critical for fast algorithm convergence: when compared with random initialization, it needs significantly fewer training iterations. We also compare different hyperparameter configurations for the same bitwidth, varying the number of codebooks and group size. Second, to validate our calibration fine-tuning procedure, we compare it against 1) no fine-tuning, 2) fine-tuning only of non-linear layers (e.g. RMSNorm) but not of codebook parameters, and 3) fine-tuning only the codebooks (but not other layers). The results, presented in full in Appendix E, show that fine-tuning the codebook parameters has the highest impact on accuracy, by far, while fine-tuning the RMSNorm only has minor impact. This validates our choice of leveraging the calibration set for learned codebooks.
Further, we observe that, increasing the number of sample sequences in the range 128 to 4096 leads to a gradual PPL improvement, but with diminishing returns. This is true for both initial AQLM calibraton and fine-tuning. In this respect, AQLM benefits more from larger calibration sets (similarly to QuIP#), as opposed to direct methods like GPTQ which saturate accuracy at around 256 input sequences. Finally, we investigate various options for investing a given bit budget, comparing e.g. longer codes (e.g. 1x15) vs multiple codebooks with shorter codes (e.g. 2x8).